Find the x-intercept and y-intercept of the line
6x-7y=10

Answer by amarjeeth123(569)

For x-intercept we have to put y=0
For y-intercept we have to put x=0
6x=10
x=10/6=5/3.
x-intercept=5/3.
-7y=10
y=-10/7
y-intercept=-10/7